# On the Localization of the Christoffel-Darboux Kernel

### v1

@@ 3 Forbidden region for the partial Christoffel-Darboux kernels
Following [18], we define the non-negligible psh envelope as
$\phi_{\mu}\mathrel{\mathop{\ordinarycolon}}=\sup\Big{\{}\phi_{0}\in{\textrm{PSH}}(X,\omega)\mathrel{\mathop{\ordinarycolon}}\phi_{0}\leq 0\text{ $\mu$-almost everywhere}\Big{\}},$ (3.2)
where ${\textrm{PSH}}(X,\omega)$ stands for the space of $\omega$-psh functions $\phi\mathrel{\mathop{\ordinarycolon}}X\to[-\infty,+\infty[$; those are upper semi-continuous functions so that $\omega+\sqrt{-1}\partial\overline{\partial}\phi$ is positive as a $(1,1)$-current. It was established by Guedj-Lu-Zeriahi [18, Proposition 2.10], cf. also [15, Proposition 3.1], that $\phi_{\mu}$ is bounded from above on $X$. To make this article self-contained, we recall briefly the argument from [15]. Following Siciak [27] and Guedj-Zeriahi [19], for a subset $E$, we define
@@
$\phi_{E}=\sup\Big{\{}\phi_{0}\in{\textrm{PSH}}(X,\omega)\mathrel{\mathop{\ordinarycolon}}\phi_{0}\leq 0\text{ over }E\Big{\}}.$ (3.3)
It is classical that if $E$ is non-pluripolar, then $\phi_{E}$ is bounded from above, cf. [19, Theorem 9.17].
Immediately from our definitions, we have $\phi_{\mu}=\sup\{\phi_{X\setminus F}\mathrel{\mathop{\ordinarycolon}}\mu(F)=0\}$. Suppose now, for contradiction, that $\phi_{\mu}$ is not bounded from above on $X$. Then there exists a sequence of sets $E_{i}$, $i\in\mathbb{N}$, with $\mu(E_{i})=0$, such that $\sup_{X}\phi_{X\setminus E_{i}}\to+\infty$, as $i\to\infty$. Then for $E\mathrel{\mathop{\ordinarycolon}}=\cup_{i\in\mathbb{N}}E_{i}$, we obviously have $\phi_{X\setminus E_{i}}\leq\phi_{X\setminus E}$ for any $i\in\mathbb{N}$, and as a consequence, $\sup_{X}\phi_{X\setminus E}=+\infty$. However, since $\mu(E)=0$, we have $\mu(X\setminus E)=1$, and so $X\setminus E$ is non-pluripolar, as we have assumed that $\mu$ is non-pluripolar. But then from [19, Theorem 9.17], $\sup_{X}\phi_{X\setminus E}<+\infty$, giving us the required contradiction.
From this point, the proof of Lemma 3.1 follows immediately from the following statement of [15, Theorem 7.1]: for any continuous function $\phi_{0}\mathrel{\mathop{\ordinarycolon}}X\to\real$, verifying $\phi_{\mu}\leq\phi_{0}$, the following holds. For any $\epsilon>0$, there is $C>0$, such that for any $k\in\mathbb{N}$, we have
${\textrm{Hilb}}_{k}(h^{L},\mu)\geq C\cdot\exp(-\epsilon k)\cdot{\textrm{Ban}}_{k}^{\infty}(h^{L}\cdot\exp(-\phi_{0})).$ (3.4)
Note that [15, Theorem 7.1] was stated for ample line bundles. However, since the analytic arguments essentially rely on [3], which applies to big line bundles, the conclusions remain valid in this more general setting. ∎
@@
Recall that Berman in [2, Lemma 4.1] established the following identity
@@
Note that in [2, Lemma 4.1], it was assumed that $E$ is irreducible, but the proof holds without this assumption. Note that our assumption on $\epsilon>0$ implies that ${\textrm{PSH}}(X,\omega_{\epsilon})$ is non-empty.
Now, let $K$ be an arbitrary non-pluripolar subset not intersecting $E$ (for example, a small ball in some local coordinates around $x\notin E$). Then for $C\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in K}(-\epsilon\cdot\sum_{i=1}^{r}a_{i}\cdot\log|s_{E_{i}}(x)|)$, we have $C<+\infty$. Let us now define $\phi_{K}\mathrel{\mathop{\ordinarycolon}}=\sup\{\phi\in{\textrm{PSH}}(X,\omega_{\epsilon})\mathrel{\mathop{\ordinarycolon}}\phi\leq C\text{ on }K\}$. The non-pluripolarity of $K$ and non-emptiness of ${\textrm{PSH}}(X,\omega_{\epsilon})$ assure that $\phi_{K}$ is uniformly bounded from above, cf. [19, Theorem 9.17]. The bound of Lemma 3.3 follows immediately from this, (3.9) and the obvious bound $\phi^{\epsilon,E}_{0}\leq\phi_{K}$. ∎
@@
Note first that it suffices to establish the result for $\epsilon>0$ small enough. From now on, we choose $\epsilon$ as in Lemma 3.3. Remark first the following characterization of the partial Christoffel-Darboux kernel
$B_{k}^{\epsilon,Y}(x,x)=\sup\frac{|s(x)|_{(h^{L})^{k}}}{\|s\|_{{\textrm{Hilb}}_{k}(h^{L},\mu)}},$ (3.11)
where the supremum is taken over all $s\in H^{0}(X,L^{\otimes k}\otimes\mathscr{O}_{X}(-\lceil\epsilon kE\rceil))\setminus\{0\}$. Now, according to Lemma 3.1, there is a continuous metric $h^{L}_{0}$ on $L$, so that for any $k\in\mathbb{N}$, we have
@@
where $\phi^{\epsilon,E}_{0}$ is defined analogously to (3.7) where we take a reference metric $h^{L}_{0}$ instead of $h^{L}$. By the definition of the $L^{\infty}$-norm, we have
$\|s\|_{{\textrm{Ban}}_{k}^{\infty}(X,h^{L}_{0}\cdot\exp(-\phi^{\epsilon,E}_{0}))}\geq|s(x)|_{(h^{L}_{0})^{k}}\cdot\exp(-k\cdot\phi^{\epsilon,E}_{0}(x))).$ (3.14)
Hence, from (3.11)-(3.14), we deduce that there is $C>0$, so that for any $x\in X$, we have
$B_{k}^{\epsilon,Y}(x,x)\leq\exp(Ck+k\cdot\phi^{\epsilon,E}_{0}(x)).$ (3.15)
The result then follows immediately from this and Lemma 3.3. ∎
4 Reduction to measures supported away from subvarieties
The main goal of this section is to show that if a given measure is truncated around a subvariety, then the associated Christoffel-Darboux kernel does not change much. As a consequence of this, we establish that in order to show Theorem 1.1 in full generality, it suffices to establish it solely for the measures supported away from a fixed subvariety.
It turns out that this form of “continuity under truncation” extends beyond subvarieties to a broader class – namely, those subsets that are not charged by any subsequential limit of the diagonal Christoffel-Darboux measures. For potential future applications, we therefore state our results in this more general setting.
More precisely, we fix a closed subset $A\subset X$ and consider a decreasing sequence of open neighborhoods $U_{1}\Supset U_{2}\Supset\cdots$ of $A$ such that $\bigcap_{i=1}^{+\infty}U_{i}=A$. Such a sequence always exists since $A$ is closed. By Urysohn’s lemma, we consider bump functions $\rho_{i}\mathrel{\mathop{\ordinarycolon}}X\to[0,1]$, $i\in\mathbb{N}^{*}$, verifying $\rho_{i}=1$ over $X\setminus U_{i}$ and $\rho_{i}=0$ over $U_{i+1}$. We assume that for any $\epsilon>0$, there is $i\in\mathbb{N}^{*}$, so that for any $k\in\mathbb{N}$, we have
$\int_{U_{i}}\mu_{k}^{\rm{Berg},\Delta}\leq\epsilon.$ (4.1)
We denote $\mu_{i}\mathrel{\mathop{\ordinarycolon}}=\rho_{i}\cdot\mu$, and let $B_{k,i}(x,y)\in L^{k}_{x}\otimes(L^{k}_{y})^{*}$, be the Christoffel-Darboux kernel associated with $h^{L}$ and $\mu_{i}$. We denote by $\mu_{k,i}^{\rm{Berg}}$ the associated probability measures on $X\times X$, defined as in (1.5). The main result of this section goes as follows.

### v2

@@ 3.2 Forbidden region for the partial Bergman kernels
For non-pluripolar measures $\mu$, Proposition 3.3 was established in [33, Theorem 7.1], where the author also provided an explicit description of $h^{L}_{0}$ in terms of the non-negligible psh envelope from [37]. An analogous explicit description can be given in our setting when $\mu$ does not give full mass to pluripolar subsets, but we decided to omit the details here for brevity.
@@
We work on the level of potentials instead of metrics. Without loosing the generality, we can assume that $h^{L}$ is smooth. We then denote $\omega\mathrel{\mathop{\ordinarycolon}}=2\pi c_{1}(L,h^{L})$. We shall establish the following result: there is $C>0$, such that for any $\psi\in{\rm{PSH}}(X,\omega)$, $p\geq 1$,
$\sup_{X}\big(\exp(p\cdot\psi)\big)\leq\exp(Cp)\cdot\int_{X}\exp(p\cdot\psi)d\mu.$ (3.11)
Once we establish (3.11), by plugging in $\psi\mathrel{\mathop{\ordinarycolon}}=\frac{1}{k}\log|s|_{h^{L^{\otimes k}}}$ and $p=2k$ in (3.11), we would get (3.10) for $h^{L}_{0}\mathrel{\mathop{\ordinarycolon}}=h^{L}\cdot\exp(-C)$.
We normalize $\mu$ so that it becomes a probability measure. For $p\in]0,+\infty[$, we define
@@
$F_{p}(\psi)\mathrel{\mathop{\ordinarycolon}}=\frac{1}{p}\log\int_{X}\exp(p\cdot\psi)d\mu,\quad F(\psi)\mathrel{\mathop{\ordinarycolon}}=\sup_{X}(\psi),\quad\text{for any }\psi\in{\rm{PSH}}(X,\omega).$ (3.12)
The generalized mean inequality implies that $F_{p}$ is a non-decreasing function of $p$. Hence, to establish (3.11), it would suffice to show that there is $C\in\real$, so that
$F_{1}(\psi)-F(\psi)\geq C,\quad\text{for any }\psi\in{\rm{PSH}}(X,\omega).$ (3.13)
Remark now that $F_{1}-F$ is clearly invariant by translation (adding a constant to the parameter), thus it descends on a function on the quotient space, which is isomorphic with the space of positive $(1,1)$-currents lying in the cohomology class of $c_{1}(L)$, which we denote by $\mathcal{T}(X,\omega)$. The latter space is compact (in the weak topology of currents), cf. [39, Proposition 8.5].
The functional $F$ is continuous by Hartogs’ lemma, cf. [39, Theorem 1.46]. Note that by [5, Lemma 1.14], the functional $F_{1}$ is also continuous if the measure $\mu$ is non-pluripolar. In this case the bound (3.13) is immediate, since continuous functions are bounded on compact sets. In the general setting, we shall adapt the proof of [5, Lemma 1.14] to obtain the desired lower bound.
Assume – for the sake of contradiction – that there exists a sequence of elements $u_{i}\in\mathcal{T}(X,\omega)$ such that $F_{1}(u_{i})-F(u_{i})\to$-$\infty$, or, equivalently,
$\int_{X}\exp(u_{i})d\mu\to 0,$ (3.14)
where – by normalizing – we view $u_{i}$ as a sequence in ${\rm{PSH}}(X,\omega)$ satisfying $F(u_{i})=0$. Recall that Hartogs’ lemma states that, up to passing to a subsequence, there exists $u\in{\rm PSH}(X,\omega)$ such that $u_{i}\to u$ in $L^{1}(X)$, and moreover
$\limsup_{i\to\infty}u_{i}=u\quad\text{outside of a pluripolar subset }E.$ (3.15)
By taking again a subsequence, we can further assume that $u_{i}$ converge to $u$ Lebesgue almost everywhere, as $i\to\infty$.
By repeating the functional-analytic argument of [5, Lemma 1.14], based only on the Banach-Alaoglu theorem, we may, for each $i\in\mathbb{N}$, find a convex combination
$v_{i}\mathrel{\mathop{\ordinarycolon}}=\sum_{j\in I_{i}}t_{i,j}\cdot\exp(u_{j}),$ (3.16)
where $I_{i}\subset\{i,i+1,\ldots\}$ is a finite subset, such that $v_{i}\to v$ in $L^{2}(\mu)$ for some $v$. Hence, upon passing to a subsequence, there exists a Borel subset $F\subset X$ with $\mu(F)=0$ such that $v_{i}\to v$ pointwise on $X\setminus F$. By (3.14) and (3.16), we then conclude that $v=0$ on $X\setminus F$.
At the same time, since $u_{i}$ converge to $u$ Lebesgue almost everywhere, we see that $v_{i}$ converge to $\exp(u)$ Lebesgue almost everywhere. By the Hartogs’ lemma, we conclude that outside of a pluripolar subset $H\subset X$, $\log(v_{i})$ converges towards $u$ (this is applicable since standard properties of plurisubharmonic functions ensure that $\log(v_{i})\in{\rm{PSH}}(X,\omega)$). We note $G\mathrel{\mathop{\ordinarycolon}}=\{x\in X\mathrel{\mathop{\ordinarycolon}}u(x)=-\infty\}$. By (3.15) and the above remark, we deduce that $X\setminus F\subset G\cup E\cup H$. Note that $G\cup E\cup H$ is a pluripolar subset verifying $\mu(G\cup E\cup H)\geq\mu(X\setminus F)=\mu(X)$, contradicting our initial assumption that $\mu$ does not give full mass to pluripolar subsets. ∎
@@
Recall that Berman in [10, Lemma 4.1] established the following identity
@@
Note that in [10, Lemma 4.1], it was assumed that $E$ is irreducible, but the proof holds without this assumption. Note that our assumption on $\epsilon>0$ implies that ${\textrm{PSH}}(X,\omega_{\epsilon})$ is non-empty.
Now, let $K$ be an arbitrary non-pluripolar subset not intersecting $E$ (for example, a small ball in some local coordinates around $x\notin E$). Then for $C\mathrel{\mathop{\ordinarycolon}}=\sup_{x\in K}(-\epsilon\cdot\sum_{i=1}^{r}a_{i}\cdot\log|s_{E_{i}}(x)|)$, we have $C<+\infty$. Let us now define $\phi_{K}\mathrel{\mathop{\ordinarycolon}}=\sup\{\phi\in{\textrm{PSH}}(X,\omega_{\epsilon})\mathrel{\mathop{\ordinarycolon}}\phi\leq C\text{ on }K\}$. The non-pluripolarity of $K$ and non-emptiness of ${\textrm{PSH}}(X,\omega_{\epsilon})$ assure that $\phi_{K}$ is uniformly bounded from above, cf. [38, Theorem 9.17]. The bound of Lemma 3.5 follows immediately from this, (3.21) and the obvious bound $\phi^{\epsilon,E}_{0}\leq\phi_{K}$. ∎
@@
Note first that it suffices to establish the result for $\epsilon>0$ small enough. From now on, we choose $\epsilon$ as in Lemma 3.5. Remark first the following characterization of the partial Bergman kernel
$B_{k}^{\epsilon,Y}(x,x)=\sup\frac{|s(x)|_{(h^{L})^{k}}^{2}}{\|s\|_{{\textrm{Hilb}}_{k}(h^{L},\mu)}^{2}},$ (3.23)
where the supremum is taken over all $s\in H^{0}(X,L^{\otimes k}\otimes\mathscr{O}_{X}(-\lceil\epsilon kE\rceil))\setminus\{0\}$. Now, according to Proposition 3.3, there is a continuous metric $h^{L}_{0}$ on $L$, so that for any $k\in\mathbb{N}$, we have
@@
where $\phi^{\epsilon,E}_{0}$ is defined analogously to (3.19) where we take a reference metric $h^{L}_{0}$ instead of $h^{L}$. Note, however, that by Lemma 3.5, there is a neighborhood $U$ of $Y$ so that
${\textrm{Ban}}_{k}^{\infty}(h^{L}_{0}\cdot\exp(-\phi^{\epsilon,E}_{0}))\geq{\textrm{Ban}}_{k}^{\infty}(U,\exp(\delta)\cdot h^{L}).$ (3.26)
By the definition of the $L^{\infty}$-norm, for any $x\in U$, we have
$\|s\|_{{\textrm{Ban}}_{k}^{\infty}(U,e\cdot h^{L})}\geq\exp(k\delta)\cdot|s(x)|_{(h^{L})^{k}}.$ (3.27)
Hence, from (3.23)-(3.27), we conclude. ∎
3.3 Diagonal Bergman measures put no mass on closed pluripolar subsets
The main result of this section refines Theorem 2.3 by showing that any subsequential limit of diagonal Bergman measures assigns no mass to pluripolar subsets. Although this result is not needed in the subsequent sections, we have included it because the proof may be of independent interest, and it highlights a natural question discussed in the end of the section.
Theorem 3.6.
For any Borel measure $\mu$ which does not give full mass to pluripolar subsets, any closed pluripolar subset $E\subset X$ and any $\epsilon>0$, there exists an open neighborhood $U$ of $E$ and $k_{0}\in\mathbb{N}$, so that for any $k\geq k_{0}$, we have
$\int_{U}\mu_{k}^{\rm{Berg},\Delta}\leq\epsilon.$ (3.28)
From Theorem 3.6, we obtain the following result.
Corollary 3.7.
Any subsequential limit of $\mu_{k}^{\rm{Berg},\Delta}$ doesn’t place mass on pluripolar subsets.
Before proceeding with the proof of Corollary 3.7, we note that Theorem 3.6 immediately implies that any subsequential limit of $\mu_{k}^{\rm{Berg},\Delta}$ doesn’t place mass on $F_{\sigma}$-pluripolar subsets, that is, to sets which can be expressed as a countable union of closed subsets. Note, however, that not every complete pluripolar subset is an $F_{\sigma}$, as illustrated by the following example kindly communicated to us by Franck Wielonsky.
Example. Let $C\subset\mathbb{C}$ be a polar Cantor set, and let $D\subset C$ be a dense countable subset. Consider the set $E\mathrel{\mathop{\ordinarycolon}}=C\setminus D$. We claim that $E$ is complete polar (i.e., there exists a subharmonic function $u$ such that $E=\{x\mathrel{\mathop{\ordinarycolon}}u(x)=-\infty\}$) and that it is not $F_{\sigma}$. First, since $C$ is closed, it is also a $G_{\delta}$ set, i.e. a countable intersection of open sets. Similarly, the complement of the countable set $D$, $\mathbb{C}\setminus D=\cap_{p\in D}(\mathbb{C}\setminus\{p\})$, is a $G_{\delta}$ set. Therefore, $E=C\cap(\mathbb{C}\setminus D)$ is also a $G_{\delta}$ set. Moreover, $E$ is polar, since it is contained in the polar set $C$. By Deny’s theorem [29] (cf. [53, §3.5]), it follows that $E$ is complete polar. To see that $E$ is not $F_{\sigma}$, suppose for contradiction that it is. Then we can write $E=\bigcup_{n=1}^{\infty}F_{n}$, where $F_{n}$ are closed. Since $D$ is dense in $C$, each $F_{n}\subset E$ is nowhere dense in $C$. Also since $C$ is a perfect set, every point of it is nowhere dense in it. Consequently, $C$ would be covered by the closed nowhere dense sets $F_{n}$ together with the points of $D$, giving a countable union of closed nowhere dense sets covering $C$. This contradicts Baire’s theorem, as $C$ is a complete metric space. Hence, $E$ cannot be $F_{\sigma}$. We also note that higher-dimensional analogues of the above phenomenon can be obtained by considering the sets $E\times\mathbb{C}^{r}\subset\mathbb{C}^{r+1}$. These sets are complete pluripolar, since $E$ is complete polar, but they are not $F_{\sigma}$, as the property of not being $F_{\sigma}$ is preserved under taking products with $\mathbb{C}^{r}$.
Proof of Corollary 3.7.
Let us consider a subsequential limit of $\mu_{k}^{\rm{Berg},\Delta}$, which we denote by $\nu$. Being a limit of Radon measures, $\nu$ is Radon. In particular, it is inner regular, see [31, Definition 1.9 and Theorems 1.8, 1.39, 1.41]. Note that any complete pluripolar subset $E$ is a $G_{\delta}$ set (in fact, if $E=\{x\in X\mathrel{\mathop{\ordinarycolon}}\psi(x)=-\infty\}$, where $\psi$ is a (local) psh function, then $E=\cap_{n\in\mathbb{N}}\{x\in X\mathrel{\mathop{\ordinarycolon}}\psi(x)<-n\}$). In particular, $E$ is Borel. By inner regularity of $\nu$, we deduce
$\nu(E)=\sup\Big\{\nu(K)\mathrel{\mathop{\ordinarycolon}}K\text{ compact},K\subset E\Big\}.$ (3.29)
Note, however, that any $K$ as above, being a subset of $E$, is also pluripolar. Hence, we have $\nu(K)=0$ by Theorem 3.6. But then by (3.29), we have $\nu(E)=0$. This shows that any subsequential limit of $\mu_{k}^{\rm{Berg},\Delta}$ doesn’t place mass on complete pluripolar subsets. As an arbitrary pluripolar subset is a subset of a complete pluripolar subsets, it finishes the proof. ∎
Despite the similarity between the statements of Theorems 2.3 and 3.6, their proofs differ drastically, as the analogue of (3.4) doesn’t hold when $\mathcal{J}_{Y}$ is replaced by the sheaf of holomorphic sections vanishing along $E$ (for instance, consider the case when $E$ is an infinite set of points with a single accumulation point). The following statement remedies this problem.